An aquarium measure 24 inches by 12 inches by 18 inches. It is filled with water to 2/3 of it's capacity (volume). What is the volume of water in the aquarium?
Found 2 solutions by Alan3354, flame8855:
Answer by Alan3354(69443)

An aquarium measure 24 inches by 12 inches by 18 inches. It is filled with water to 2/3 of it's capacity (volume). What is the volume of water in the aquarium?
-------------
The volume is the product of the 3 dimensions in cubic inches.
------
The water in it is 2/3 of that.
----
it's = it is

Answer by flame8855(424)

volume = 24*12*18 = 5184 in^3
2/3 volume = 3456 volume of water